OpenStudy (anonymous):

What are the factors of x2 + 11x + 24? (x − 6)(x − 4) (x + 6)(x + 4) (x − 3)(x − 8) (x + 3)(x + 8)

OpenStudy (anonymous):

can u explain ?

OpenStudy (e.cociuba):

My bad hun, it's C.) x^2 - 11x + 24 { x^2 means x squared } = x^2 - 8x - 3x + 24............{ Since, 24 = (- 8) x (- 3) and - 11x = - 8x - 3x } = x (x - 8) - 3 (x - 8) = (x - 3)(x - 8)

OpenStudy (e.cociuba):

Wat do ya'll say @skullpatrol

OpenStudy (anonymous):

*y'all looks good to me :)

OpenStudy (e.cociuba):

Haha Oops srry :) KK :)

OpenStudy (anonymous):

how about this one = What are the factors of 3x2 + 10x + 3? (3x − 1)(x − 3) (3x − 3)(x − 1) (3x + 1)(x + 3) (3x + 3)(x + 1)

OpenStudy (e.cociuba):

x+3 and 3x+1 are both factors of the polynomial given. So, it's C.)

OpenStudy (e.cociuba):

3x^2+10^x+3 =3x^2+9x+x+3[by middle term factorization method] =3x(x+3)+(x+3) =(x+3)(3x+1) U can use F.O.I.L as proof.....

OpenStudy (skullpatrol):

FOIL is what I showed you on the last question.

OpenStudy (e.cociuba):

who??

OpenStudy (anonymous):

me. good job

OpenStudy (e.cociuba):

Do u get it??